OpenStudy (anonymous):

In the following example of "Simplifying an Expression", what reason would be used for step A? Simplify the expression 3k + 7 + 2k and justify each step 3k + 7 + 2k = 3k + 2k + 7 A = (3k+2k) +7 B = (3+2)k+7 C = 5k + 7 D A. Addition B. Commutative Property C. Associative Property D. Distributive Property

OpenStudy (anonymous):

Oct 29, 2010 - Simplify the expression 3k + 7 + 2k and justify each step 3k + 7 + 2k = 3k + 2k + 7 A = (3k + 2k) + 7 B = (3 + 2)k + 7 C = 5k +